OpenStudy (anonymous):

Find the exact value of sin 375

OpenStudy (unklerhaukus):

375 = 360+15

OpenStudy (unklerhaukus):

15 = 45-30

OpenStudy (unklerhaukus):

sin(a-b) = sin(a) cos(b) - cos(a) sin(b)

Nnesha (nnesha):

there are 3+ options to find exact value you need to break 375 into two angles that are on the unit circle here is another (150+225= 375 )

OpenStudy (anonymous):

Oh ok I get it! Thanks!